OpenStudy (anonymous):

What is the sum of this infinite geometric series? ∞ ∑(12)(0.5)^i i=1

OpenStudy (campbell_st):

well list a few terms i = 1 6 i = 2 3 i = 3 1.5 i = 4 0.75 to the common ratio r = 1/2 the 1st term is 6 so the infinite sum can be found by \[S_{\infty} = \frac{a}{(1 - r)}\]